# Gröbner Basis Cryptanalysis of Ciminion and Hydra

### v1

@@ 4.1 Generic Coordinates for Hydra
$\mathbf{g}_{j}^{(i)}$ $=\left(\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{j}^{(i)}\right)^{\text{hom}}\mod x_{0},$
$\mathbf{g}_{\mathcal{R}}$ $=\left(\mathbf{B}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{f}_{\mathcal{R}}\right)^{\text{hom}}\mod x_{0}.$
@@
$\mathbf{g}_{j}^{(i)}=\begin{pmatrix}\left\{x_{j,k}^{(i)}-x_{j,8}^{(i)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{j}^{(i+1)}\Big{)}\bigg{)}_{k}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{j}^{(i+1)}-\mathbf{k}\Big{)}\bigg{)}_{8}\right\}_{1\leq k\leq 7}\\[10.0pt] \left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{j,l}^{(i)}\right)^{2}\end{pmatrix}.$
@@
$\mathbf{B}\mathbf{g}_{\mathcal{R}}=\begin{pmatrix}\left\{y_{k}-y_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{k}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{4}\right\}_{1\leq k\leq 3}\\[10.0pt] g(\mathbf{y},\mathbf{z})\\ \left\{z_{k}-z_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{4+k}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{8}\right\}_{1\leq k\leq 3}\\[10.0pt] g(\mathbf{z},\mathbf{y})\end{pmatrix}.$
Except for the rolling function, all quadratic functions are powers of affine equations, so by Theorem 2.6 (2) we can replace the powers by the affine equations. This yields the linear system from the assertion. ∎
Naturally, this theorem can be extended to an arbitrary number of Hydra samples. It also follows by a simple contraposition that the downsized Hydra polynomial system in $42\cdot r_{\mathcal{H}}+$ 4 variables is in generic coordinates if the full polynomial system is in generic coordinates. Now it is straight-forward to compute the Macaulay bound (Theorem 2.5) for Hydra.
@@
Let $\mathbb{F}_{q}$ be a finite field, let $\overline{\mathbb{F}_{q}}$ be its algebraic closure, and let
~$\widetilde{\mathcal{F}}_{\textsf{Hydra}}\subset\overline{\mathbb{F}_{q}}\Big{[}z_{4},x_{1,8}^{(1)},\dots,x_{1,8}^{(r_{\mathcal{H}}-1)},x_{2,8}^{(0)},\dots,x_{2,8}^{(r_{\mathcal{H}}-1)},\mathbf{k}\Big{]}$
be the downsized Hydra polynomial system. If ~$\widetilde{\mathcal{F}}_{\textsf{Hydra}}^{\text{hom}}$ is in generic coordinates, then
$\operatorname{sd}_{DRL}\left(\widetilde{\mathcal{F}}_{\textsf{Hydra}}\right)\leq 2\cdot r_{\mathcal{H}}+3.$
Remark 4.3.
This bound can be further improved by utilizing the $6$ affine equations from the rolling function. In case that these $6$ equations are linearly independent of each other and all other affine equations, we can eliminate further $6$ variables. Thus, the Hydra polynomial system reduces to $22\cdot r_{\mathcal{H}}+$ 2 quadratic polynomials in $22\cdot r_{\mathcal{H}}-$ 2 many variables. Then
$\operatorname{sd}_{DRL}\left(\widetilde{\mathcal{F}}_{\textsf{Hydra}}\right)\leq 2\cdot r_{\mathcal{H}}.$
The reason why we did not incorporate this improvement in Theorem 2.5 is that elimination via the affine equations from the rolling function depends non-trivially on $\mathbf{M}_{\mathcal{R}}$ and $\mathbf{M}_{\mathcal{J}}$. Thus, a generic treatment is not possible.
@@ 4.2 Extracting a Zero-Dimensional DRL Gröbner Basis for Hydra
Under the assumption of Remark 4.3, we can extract a zero-dimensional Gröbner basis from the downsized Hydra polynomial system via a linear change of coordinates. Then, $2\cdot r_{\mathcal{H}}$ quadratic polynomials stem from Hydra heads and the remaining two from the rolling function.
Recall that the quadratic terms in the heads are of the form
$g_{1}^{(1)}$ $=\left(\sum_{l=1}^{4}y_{l}-z_{l}\right)^{2},\qquad$ $g_{2}^{(1)}=\left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{2,l}^{(0)}\right)^{2},$ (40)
$g_{j}^{(i)}$ $=\left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{j,l}^{(i)}\right)^{2},$ $2\leq i\leq r_{\mathcal{H}}-1,\ j\in\{1,2\}$ (41)
@@
$g_{1}^{(r_{\mathcal{H}})}$ $=\left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{1,l}^{(r_{\mathcal{H}}-1)}\right)^{2},\qquad$ $g_{2}^{(r_{\mathcal{H}})}=\left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{2,l}^{(r_{\mathcal{H}}-1)}\right)^{2},$ (42)
i.e. they are the squares of linear combinations of the variables. Therefore, we collect these linear combinations in a matrix $\mathbf{M}\in\mathbb{F}_{q}^{2\cdot r_{\mathcal{H}}\times(2\cdot r_{\mathcal{H}}-2)}$ analog to the linear system in Theorem 4.1 to verify generic coordinates for Hydra. If $\operatorname{rank}\left(\mathbf{M}\right)=2\cdot r_{\mathcal{H}}-2$, then we can perform a linear change of coordinates to new variables $\hat{x}_{1},\dots,\hat{x}_{2\cdot r_{\mathcal{H}}-2}$. After the change of coordinates the Hydra polynomial system contains $22\cdot r_{\mathcal{H}}-$ 2 polynomials of the form
$g_{i}=\hat{x}_{i}^{2}+\sum_{j=1}^{2\cdot r_{\mathcal{H}}-2}\alpha_{i,j}\cdot\hat{x}_{j}+\alpha_{i,0}.$ (43)
Trivially, $\mathcal{G}_{\textsf{Hydra}}=\{g_{i}\}_{1\leq i\leq 2\cdot r_{\mathcal{H}}-2}\subset\mathbb{F}_{q}\big{[}\hat{x}_{1},\dots,\hat{x}_{2\cdot r_{\mathcal{H}}-2}\big{]}$ forms a zero-dimensional DRL Gröbner basis, see [CLO15, Chapter 2 §9 Theorem 3, Proposition 4]. In particular, for the $\mathbb{F}_{q}$-vector space dimension of the quotient ring we then have that
@@
$\dim_{\mathbb{F}_{q}}\left(\mathcal{G}_{\textsf{Hydra}}\right)=2^{2\cdot r_{\mathcal{H}}-2}.$ (44)
We also note that after an auxiliary division by remainder computation, the additional four polynomials are of the form
$h_{j}=\sum_{1\leq k<l\leq 2\cdot r_{\mathcal{H}}-2}\alpha_{i,k,l}\cdot\hat{x}_{k}\cdot\hat{x}_{l}+\sum_{k=1}^{2\cdot r_{\mathcal{H}}-2}\beta_{i,k}\cdot\hat{x}_{k}+\beta_{i,0}.$ (45)
@@ 4.3 Implementation
We have implemented Hydra and its iterated polynomial system in the computer algebra system SageMath [Sag24] and OSCAR [OSC24]. [^4] https://github.com/sca-research/Groebner-Basis-Cryptanalysis-of-Ciminion-and-Hydra.git In principle, the SageMath implementation can handle $128$ bit prime numbers, though we do not recommend using primes of that size for experimental Gröbner basis computations, because SageMath uses Python implementations for primes of that size. The OSCAR computer algebra system offers an efficient F4 implementation that can handle primes of size $q<2^{31}$, therefore we recommend to use OSCAR for experimental computations over small prime fields.
Moreover, we implemented the linear system from Theorem 4.1 in SageMath to verify whether a Hydra instance is in generic coordinates. We stress that our generic coordinates test can handle practical parameters like $128$ bit prime numbers. Also, our verification function can handle an arbitrary number of Hydra samples.
In case a Hydra polynomial system is in generic coordinates, we implemented the change of coordinates from Section 4.2 for two samples to extract a zero-dimensional DRL Gröbner basis from the Hydra polynomial system.
4.3.1 A Concrete Instance
@@
$\mathbf{M}_{\mathcal{E}}$ $=\operatorname{circ}\left(3,2,1,1\right),$ (46)
$\mathbf{M}_{\mathcal{I}}$ $=\begin{pmatrix}1&1&1&1\\ 1&4&1&1\\ 3&1&3&1\\ 4&1&1&2\end{pmatrix},$ (47)
@@
For security level $\kappa=128$ the designers suggest to use $r_{\mathcal{H}}=39$, see [GØSW23, §5.5].
@@
For this instance using our SageMath implementation we verified that the Hydra polynomial system is in generic coordinates for $m\in\{2,3,4,5,10\}$ many samples. For $m=2$ we can compute a Hydra DRL Gröbner basis with the change of coordinates from Section 4.2. In particular, the polynomial system then consists of $22\cdot r_{\mathcal{H}}+$ 2 polynomials in $22\cdot r_{\mathcal{H}}-$ 2 variables, so we can use the slightly improved solving degree bound from Remark 4.3 to estimate the complexity of DRL Gröbner basis computations. On an Intel Core i7-10700 with 64 GB RAM verification of generic coordinates for $m=2$ needs $\leq 10\,\text{s}$ and for $m=10$ it needs $\leq 4\,\text{min}$.
In addition, for $r_{\mathcal{H}}=39$, to extract a Gröbner basis let $\mathbf{M}\in\mathbb{F}_{q}^{2\cdot r_{\mathcal{H}}\times(2\cdot r_{\mathcal{H}}-2)}$ collect the linear combinations in the quadratic terms of the downsized Hydra polynomial system. Then, the rows with indices $3,\dots,2\cdot r_{\mathcal{H}}$ are linearly independent. We also note that we did not find a round number $r_{\mathcal{H}}\geq 2$ for which the Gröbner basis extraction fails.
5 Shaving-Off Bits with the Eigenvalue Method
Let =$\mathcal{F}=\{f_{1},\dots,f_{n}\}\subset\mathbb{F}_{q}[x_{1},\dots,x_{n}]$ be such that $\operatorname{LM}_{DRL}\left(f_{i}\right)=x_{i}^{2}$ for all $i$, then $\mathcal{F}$ is a zero-dimensional Gröbner basis. Moreover, for Ciminion and Hydra we found Gröbner bases that satisfy this assumption. In Section 2.5 we already discussed that one can solve for the zero-locus of $\mathcal{F}$ by computing the eigenvalues of the multiplication matrices while maintaining the complexity estimate of the probabilistic FGLM algorithm [FGHR14]. We will now discuss that we can slightly improve the complexity estimate by taking the multiplicative structure of $\mathbb{F}_{q}[x_{1},\dots,x_{n}]/\left(\mathcal{F}\right)$ into account.
Recall that a $\mathbb{F}_{q}$-vector space basis of $\mathbb{F}_{q}[x_{1},\dots,x_{n}]/\left(\mathcal{F}\right)$ is given by the monomials not contained in $\operatorname{in}_{DRL}\left(\mathcal{F}\right)=\left(x_{1}^{2},\dots,x_{n}^{2}\right)$. Let $\mathcal{B}^{\prime}$ be the monomials not contained in $\left(x_{1}^{2},\dots,x_{n-1}^{2}\right)\subset\mathbb{F}_{q}[x_{1},\dots,x_{n-1}]$, then we sort the quotient space basis as =$\mathcal{B}=\left\{\mathcal{B}^{\prime},x_{n}\cdot\mathcal{B}^{\prime}\right\}$. The multiplication matrix for $x_{n}$ is then of the form
[…]
where $\mathbf{A},\mathbf{B}\in\mathbb{F}_{q}^{2^{n-1}\times 2^{n-1}}$. By definition, the characteristic polynomial is given by the determinant
$\chi_{x_{n}}(x)=\det\left(\mathbf{I}_{2^{n}\times 2^{n}}\cdot x-\mathbf{M}_{x_{n}}\right)=\det\begin{pmatrix}\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x&-\mathbf{I}_{2^{n-1}\times 2^{n-1}}\\ \mathbf{A}&\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x-\mathbf{B}\end{pmatrix}.$ (50)
For the LU-decomposition from Equation 24, we lift the determinant computation from the polynomial ring $\mathbb{F}_{q}[x]$ to its fraction field $\mathbb{F}_{q}(x)$, i.e. we treat $x$ as invertible. Then with Equation 25
$\chi_{x_{n}}(x)$ $=\det\begin{pmatrix}\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x&-\mathbf{I}_{2^{n-1}\times 2^{n-1}}\\ -\mathbf{A}&\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x-\mathbf{B}\end{pmatrix}$ (51)
$=\det\left(\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x\right)\cdot\det\left(\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x-\mathbf{B}-\mathbf{A}\cdot x^{-1}\right)$ (52)
$=x^{2^{n-1}}\cdot\det\left(\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x-\mathbf{B}-\mathbf{A}\cdot x^{-1}\right)$ (53)
$=\det\left(\mathbf{I}_{2^{n-1}\times 2^{n-1}}\cdot x^{2}-\mathbf{B}\cdot x-\mathbf{A}\right),$ (54)
where the latter equality follows from the fact that the determinant is a homogeneous function. We note that a similar determinant trick has been applied in [BBL+24, §3.2] to improve characteristic polynomial computations for specially shaped Gröbner bases.
For the next variable $x_{n-1}$, we extract a $\mathbb{F}_{q}$-valued solution for $x_{n}$ via the GCD method, substitute it into $\mathcal{F}\setminus\left\{f_{n}\right\}$, and repeat the trick for the multiplication matrix, see Equation 49, with one variable less.
Let $D(d,q)$ denote the complexity of the GCD with the field equation, see Equation 1. Provided that every characteristic polynomial has at most $N$ roots in $\mathbb{F}_{q}$, then the complexity of the Eigenvalue method with standard determinant complexity is given by
$\mathcal{O}\Bigg{(}\sum_{i=1}^{n}N^{n-i}\cdot\Big{(}2^{\omega\cdot i}+D\big{(}2^{i},q\big{)}\Big{)}\Bigg{)}=\mathcal{O}\Bigg{(}2^{\omega}\cdot\frac{2^{\omega\cdot n}-N^{n}}{2^{\omega}-N}+\sum_{i=1}^{n}N^{n-i}\cdot D\big{(}2^{i},q\big{)}\bigg{)}\Bigg{)}.$ (55)
With the determinant trick from Equation 54 the complexity for the Eigenvalue Method then improves to
$\mathcal{O}\Bigg{(}\sum_{i=1}^{n}N^{n-i}\cdot\Big{(}2^{\omega\cdot(i-1)}+D\big{(}2^{i},q\big{)}\Big{)}\Bigg{)}=\mathcal{O}\Bigg{(}\frac{2^{\omega\cdot n}-N^{n}}{2^{\omega}-N}+\sum_{i=1}^{n}N^{n-i}\cdot D\big{(}2^{i},q\big{)}\bigg{)}\Bigg{)}.$ (56)
In addition, for $N=1$, this implies by $\omega\geq 2$ that
$\log_{2}\left(\frac{2^{\omega\cdot n}-1}{2^{\omega}-1}\right)\leq\omega\cdot n-\log_{2}\left(3\right).$ (57)
I.e., we expect an improvement of at least $\log_{2}\left(3\right)\approx 1.5849$ compared to generic probabilistic term order conversion for an ideal in shape position.
By imposing more structure on $\mathcal{F}$ we can refine Equation 56 even further. Suppose that every $f_{i}\in\mathcal{F}$ is of the form
$f_{i}=x_{i}^{2}+\mathcal{A}_{i}(x_{1},\dots,x_{n}),$ (58)
where $\mathcal{A}_{i}$ is an affine polynomial. Assume we have found a solution for $x_{n}$, say $x_{n}=\alpha_{n}$, then we can rearrange

### v4

@@ 4.1 Generic Coordinates for Hydra
$\mathbf{g}_{j}^{(i)}$ $=\left(\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{j}^{(i)}\right)^{\text{hom}}\mod(x_{0}),$
$\mathbf{g}_{\mathcal{R}}$ $=\left(\mathbf{B}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{f}_{\mathcal{R}}\right)^{\text{hom}}\mod(x_{0}).$
@@
$\mathbf{g}_{j}^{(i)}=\begin{pmatrix}\left\{x_{j,k}^{(i-1)}-x_{j,8}^{(i-1)}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{k}^{\prime}-\mathbf{x}_{j}^{(i)}\Big{)}\bigg{)}_{k}+\bigg{(}\mathbf{M}_{\mathcal{J}}^{-1}\Big{(}\mathbf{x}_{j}^{(i)}-\mathbf{k}^{\prime}\Big{)}\bigg{)}_{8}\right\}_{1\leq k\leq 7}\\[10.0pt] \left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{j,l}^{(i-1)}\right)^{2}\end{pmatrix}.$
@@
$\mathbf{g}_{\mathcal{R}}=\begin{pmatrix}\left\{y_{k}-y_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{k}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{4}\right\}_{1\leq k\leq 3}\\[10.0pt] g(\mathbf{y},\mathbf{z})\\ \left\{z_{k}-z_{4}-\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{4+k}+\Big{(}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{x}_{2}^{(0)}\Big{)}_{8}\right\}_{1\leq k\leq 3}\\[10.0pt] g(\mathbf{z},\mathbf{y})\end{pmatrix}.$
Except for the rolling function, all quadratic polynomials are squares of linear polynomials. Collecting the linear polynomials from $\mathcal{G}_{\textsf{Hydra}}^{\text{top}}$ together with the linear polynomials from the squares, we yield the linear system from the assertion. It is clear that all these linear polynomials are elements of $\sqrt{\mathcal{G}_{\textsf{Hydra}}^{\text{top}}}$. Since the highest degree components are homogeneous, we always have that $\sqrt{\mathcal{G}_{\textsf{Hydra}}^{\text{top}}}\subset\left(\mathbf{y},\mathbf{z},\mathbf{x}_{1}^{(i)},\dots,\mathbf{x}_{1}^{(r_{\mathcal{H}}-1)},\mathbf{x}_{2}^{(0)},\dots,\mathbf{x}_{2}^{(r_{\mathcal{H}}-1)},\mathbf{k}\right)$. Thus, if the linear system from the assertion has full rank, then we have established the equality from Equation 47. ∎
Naturally, this theorem can be extended to an arbitrary number of Hydra samples. Now it is straight-forward to compute the Macaulay bound (Theorem 2.6) for Hydra.
@@
$\mathcal{F}_{\textsf{Hydra}}=\left\{\mathbf{f}_{1}^{(i)},\mathbf{f}_{\mathcal{R}},\mathbf{f}_{2}^{(i)}\right\}_{1\leq i\leq r_{\mathcal{H}}}\subset P=\overline{\mathbb{F}_{q}}\Big{[}\mathbf{y},\mathbf{z},\mathbf{x}_{1}^{(1)},\dots,\mathbf{x}_{1}^{(r_{\mathcal{H}}-1)},\mathbf{x}_{2}^{(0)},\dots,\mathbf{x}_{2}^{(r_{\mathcal{H}}-1)},\mathbf{k}\Big{]}$
be a Hydra polynomial system. Let
$\mathcal{G}_{\textsf{Hydra}}=\left\{\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{1}^{(i)},\mathbf{B}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{f}_{\mathcal{R}},\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{2}^{(i)}\right\}_{1\leq i\leq r_{\mathcal{H}}},$
and assume that $\left(\mathcal{G}_{\textsf{Hydra}}^{\text{hom}}\right)$ is in generic coordinates. If the affine polynomials in $\mathcal{G}_{\textsf{Hydra}}$ have rank $62\cdot 7\cdot r_{\mathcal{H}}+$ 6, then
$\operatorname{sd}_{DRL}\left(\mathcal{G}_{\textsf{Hydra}}\right)\leq 2\cdot r_{\mathcal{H}}.$
Proof.
If the affine polynomials have rank $62\cdot 7\cdot r_{\mathcal{H}}+$ 6, then we can reduce to $22\cdot r_{\mathcal{H}}+$ 2 quadratic polynomials in $22\cdot r_{\mathcal{H}}-$ 2 variables. This variable elimination does not affect being in generic coordinates, so the Macaulay bound (Theorem 2.6) evaluates to
$\operatorname{sd}_{DRL}\left(\mathcal{G}_{\textsf{Hydra}}\right)\leq 2\cdot(2\cdot r_{\mathcal{H}}-1)-\left(2\cdot r_{\mathcal{H}}-1\right)+1=2\cdot r_{\mathcal{H}}.\qed$
@@ 4.2 Extracting a Zero-Dimensional DRL Gröbner Basis for Hydra
Due to the structure of the Hydra heads we can actually produce a DRL Gröbner basis via a simple linear change of coordinates. Again, let $\mathcal{G}_{\textsf{Hydra}}$ be as in Theorem 4.1, we assume that
  1. (i)
    the linear system from Theorem 4.1 has full rank, and
  2. (ii)
    the linear polynomials in $\mathcal{G}_{\textsf{Hydra}}^{\text{top}}$ have rank $62\cdot 7\cdot r_{\mathcal{H}}+$ 6.
I.e., $\left(\mathcal{G}_{\textsf{Hydra}}^{\text{hom}}\right)$ is in generic coordinates, and we can use the linear polynomials in $\mathcal{G}_{\textsf{Hydra}}$ to reduce the system to $22\cdot r_{\mathcal{H}}+$ 2 quadratic polynomials in $22\cdot r_{\mathcal{H}}-$ 2 variables.
By construction $2\cdot r_{\mathcal{H}}$ of these quadratic polynomials come from the Hydra heads and their quadratic terms are squares of linear equations. In particular, before the variable elimination the quadratic terms are, see Definition 2.2 and the proof of Theorem 4.1,
$g_{1,8}^{(1)}$ $=\left(\sum_{l=1}^{4}y_{l}-z_{l}\right)^{2}$ $=\mathcal{L}_{1,1}^{2},$ (48)
$g_{2,8}^{(1)}$ $=\left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{2,l}^{(0)}\right)^{2}$ $=\mathcal{L}_{2,1}^{2},$ (49)
@@
$g_{j,8}^{(i)}$ $=\left(\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{j,l}^{(i-1)}\right)^{2}$ $=\mathcal{L}_{j,i}^{2},\qquad 2\leq i\leq r_{\mathcal{H}},\ j\in\{1,2\},$ (50)
Recall that these linear equations =$\mathcal{L}=\{\mathcal{L}_{1,i},\mathcal{L}_{2,i}\}_{1\leq i\leq 2\cdot r_{\mathcal{H}}}$ also appeared in the linear system from Theorem 4.1. By Assumption (i), $\mathcal{L}$ together with the linear equations from $\mathcal{G}_{\textsf{Hydra}}^{\text{top}}$ has rank $416\cdot r_{\mathcal{H}}+$ 4. After eliminating $62\cdot 7\cdot r_{\mathcal{H}}+$ 6 variables with the affine equations from $\mathcal{G}_{\textsf{Hydra}}$, the polynomials from Equations 48, 49 and 50 modify to
$g_{j,8}^{(i)}=\left(\tilde{\mathcal{L}}_{j,i}+a_{j,i}\right)^{2},\qquad 1\leq i\leq r_{\mathcal{H}},\ j\in\{1,2\},$ (51)
where ~$\tilde{\mathcal{L}}_{j,i}$ is a linear polynomial in the remaining $22\cdot r_{\mathcal{H}}-$ 2 variables and $a_{j,i}\in\mathbb{F}_{q}$. Moreover, the rank of a matrix does not depend on the order of Gaussian elimination, so by Assumption (i) the new linear polynomials ~={~$\tilde{\mathcal{L}}=\left\{\tilde{\mathcal{L}}_{1,i},\tilde{\mathcal{L}}_{2,i}\right\}_{1\leq i\leq 2\cdot r_{\mathcal{H}}}$ still must contain a maximal linearly independent subset of rank $22\cdot r_{\mathcal{H}}-$ 2.
Let $\mathbf{x}_{\text{nl}}$ denote the remaining $22\cdot r_{\mathcal{H}}-$ 2 non-linear variables. We can now collect this maximal linearly independent subset of $\tilde{\mathcal{L}}$ in a matrix $\mathbf{M}\in\mathbb{F}_{q}^{(2\cdot r_{\mathcal{H}}-2)\times(2\cdot r_{\mathcal{H}}-2)}$. Columns of $\mathbf{M}$ are indexed by the elements of $\mathbf{x}_{\text{nl}}$, and in the rows of $\mathbf{M}$ we write the coefficients of the maximal linearly independent equations. Next we pick new variables $\mathbf{\hat{x}}=\left(\hat{x}_{1},\dots,\hat{x}_{2\cdot r_{\mathcal{H}}-2}\right)^{\intercal}$, and finally we perform a linear change of variables $\mathbf{\hat{x}}=\mathbf{M}\mathbf{x}_{\text{nl}}$. By construction $2\cdot r_{\mathcal{H}}$ polynomials in $\mathcal{G}_{\textsf{Hydra}}$ are of the form
~$\tilde{\mathcal{L}}_{j,i}(\mathbf{x}_{\text{nl}})^{2}+\mathcal{A}_{j,i}(\mathbf{x}_{\text{nl}})=0,$ (52)
where $\mathcal{A}_{j,i}\in\mathbb{F}_{q}[\mathbf{x}_{\text{nl}}]$ is affine, $1\leq i\leq 2\cdot r_{\mathcal{H}}$ and $j\in\{1,2\}$. Therefore, after the change of coordinates we produce $22\cdot r_{\mathcal{H}}-$ 2 polynomials of the form
$g_{i}=\hat{x}_{i}^{2}+\sum_{j=1}^{2\cdot r_{\mathcal{H}}-2}\alpha_{i,j}\cdot\hat{x}_{j}+\alpha_{i,0},$ (53)
where $\alpha_{i,j},\alpha_{i,0}\in\mathbb{F}_{q}$, together with four additional quadratic polynomials. In particular, $\hat{\mathcal{G}}=\left\{g_{i}\right\}_{1\leq i\leq 2\cdot r_{\mathcal{H}}-2}\subset\mathbb{F}_{q}\big{[}\mathbf{\hat{x}}\big{]}$ is a DRL Gröbner basis with $\dim_{\mathbb{F}_{q}}\left(\hat{\mathcal{G}}\right)=2^{2\cdot r_{\mathcal{H}}-2}$.
We summarize this construction in a more formal manner in the next corollary.
Corollary 4.3.
Let $\mathbb{F}_{q}$ be a finite field, let
@@
$\mathcal{F}_{\textsf{Hydra}}=\left\{\mathbf{f}_{1}^{(i)},\mathbf{f}_{\mathcal{R}},\mathbf{f}_{2}^{(i)}\right\}_{1\leq i\leq r_{\mathcal{H}}}\subset P=\mathbb{F}_{q}\Big{[}\mathbf{y},\mathbf{z},\mathbf{x}_{1}^{(1)},\dots,\mathbf{x}_{1}^{(r_{\mathcal{H}}-1)},\mathbf{x}_{2}^{(0)},\dots,\mathbf{x}_{2}^{(r_{\mathcal{H}}-1)},\mathbf{k}\Big{]}$
be a Hydra polynomial system, and let
$\mathcal{G}_{\textsf{Hydra}}$ =$=\left\{\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{1}^{(i)},\mathbf{B}\mathbf{M}_{\mathcal{R}}^{-1}\mathbf{f}_{\mathcal{R}},\mathbf{A}_{8}\mathbf{M}_{\mathcal{J}}^{-1}\mathbf{f}_{2}^{(i)}\right\}_{1\leq i\leq r_{\mathcal{H}}},$
$\mathcal{G}_{lin}$ =$=\left\{g\in\mathcal{G}_{\textsf{Hydra}}\mid\deg\left(g\right)=1\right\},$
$\mathcal{G}_{squ}$ =$=\left\{g\in\mathcal{G}_{\textsf{Hydra}}\mid\deg\left(g\right)=2\right\}.$
Assume that
  1. (i)
    the linear system from Theorem 4.1 has full rank, and
  2. (ii)
    $\operatorname{rank}\left(\mathcal{G}_{lin}^{\text{top}}\right)=2\cdot 7\cdot r_{\mathcal{H}}+6$.
[…]
  1. (1)
[…]
  2. (2)
    For every $21\leq i\leq 2\cdot r_{\mathcal{H}}-$ 2 there exists $g_{i}\in\hat{\mathcal{G}}_{squ}$ such that
    $g_{i}=\hat{x}_{i}^{2}+\mathcal{A}_{i},$
[…]
  3. (3)
    $\dim_{\mathbb{F}_{q}}\left(\mathcal{G}_{lin}\cup\mathcal{G}_{sub}\cup\hat{\mathcal{G}}_{squ}\right)\leq 2^{2\cdot r_{\mathcal{H}}-2}$.
Besides being tedious to do by hand, success of the change of coordinates depends on the matrices used in the Hydra heads. Therefore, we did not attempt a full proof by hand, instead for specific instances we outsourced it to a computer algebra program where the construction can be performed in a matter of seconds.
@@ 4.3 Implementation
We have implemented Hydra and its iterated polynomial system in the computer algebra systems SageMath [Sag24] and OSCAR [OSC24]. [^6] https://github.com/sca-research/Groebner-Basis-Cryptanalysis-of-Ciminion-and-Hydra.git In both systems we implemented the linear system from Theorem 4.1 and the change of coordinates from Section 4.2. With the SageMath implementation we can compute the DRL Gröbner basis for full rounds $r_{\mathcal{H}}=39$ over $128$ bit prime fields in negligible time. The OSCAR computer algebra system offers an efficient F4 implementation that can handle primes of size $q<2^{31}$, therefore we recommend to use OSCAR for experimental computations over small prime fields.
For completeness, we mention that we implemented the rolling function without a constant addition, i.e. we set $\mathbf{c}_{\mathcal{R}}^{(i)}=\mathbf{0}$ for all $i\geq 1$. However, we emphasize that the constant addition does not affect either the linear system from Theorem 4.1 or the change of coordinates from Section 4.2.
Example 4.4 (A Concrete Instance).
@@
$\mathbf{M}_{\mathcal{E}}$ $=\operatorname{circ}\left(3,2,1,1\right),\qquad\qquad\mathbf{M}_{\mathcal{I}}=\begin{pmatrix}1&1&1&1\\ 1&4&1&1\\ 3&1&3&1\\ 4&1&1&2\end{pmatrix},$ (54)
@@
For security level $\kappa=128$ bits the designers suggest to use $r_{\mathcal{H}}=39$, see [GØSW23, §5.5].
@@
For this instance using our SageMath implementation we verified that the Hydra polynomial system is in generic coordinates for $m\in\{2,3,4,5,10\}$ many samples. For $m=2$ we can compute a Hydra DRL Gröbner basis with the change of coordinates from Section 4.2.
Recall the quadratic terms from the Hydra heads, see Equations 48, 49 and 50. For $r_{\mathcal{H}}\geq 2$ and after the elimination of $62\cdot 7\cdot r_{\mathcal{H}}+$ 6 variables, see Equation 51, we observed that